OpenStudy (anonymous):

I seriously do not understand equations with variables on both sides, is there an easier or simpler way of explaning it?

hero (hero):

You mean you don't understand 5x + 6 = 3x

OpenStudy (anonymous):

no i do not, it seems to be my one weakness in math

hero (hero):

And then you solve by placing like terms on the same side... 5x - 3x + 6 = 0 5x - 3x = 0 - 6 2x = -6 x = -6/2 x = -3

OpenStudy (anonymous):

thank you, this helps

OpenStudy (anonymous):

It depends on the equation, but many are of the form y = f(x) = "some expression in x". On the right side, x will have a domain or allowable values. For simplicity, let's not restrict x and let it be any real number for now. We select an x, any x, and for that SPECIFIC chosen x we will perform operations on it and get a "y". Since we start with "x" and the value of "y" is dependent on the "x" selected, we call "x" the independent variable and "y" the dependent variable. Is this helping or is your confusion in a different area of variables?

OpenStudy (anonymous):

well one i don't understand is -1/2x + 10 = −x + 12

hero (hero):

You do the same thing bro.

hero (hero):

Place like terms on the same side

OpenStudy (anonymous):

One fundamental consideration is that "ax" means a times x. Also, ax + bx = (a + b)x. It is necessary to understand at least this much before moving variables from one side to the other.

hero (hero):

-.5x + 10 = - x + 12 x - .5x = 12 - 10 .5x = 2 x = 2/.5 x = 4

hero (hero):

I change 1/2 to .5 because they mean the same thing

OpenStudy (anonymous):

okay then thank you